# A Literature Review of Literature Reviews in Pattern Analysis and Machine Intelligence

### v5

@@ 7.1 Challenges
Lag and Obsolescence.
Scientific knowledge is constantly evolving, yet the process of compiling reviews often struggles to keep up with the latest research advancements—particularly when created by human authors and subject to peer review. Comprehensive reviews are often delayed relative to the emergence of new research areas. For instance, when Mamba [34] started gaining traction, no in-depth reviews were available during its initial months. This delay reflects the time-intensive nature of gathering, evaluating, and synthesizing an ever-expanding body of literature, along with the requirements of the peer-review process.
@@ 7.2 Future
AI Empowerment Literature Review. While debates persist regarding the ethical implications and appropriateness of AI-generated content for formal academic publication, such AI-driven reviews have gained acceptance for informal purposes, e.g., rapid ad-hoc reviews [101]. Considering the substantial time and effort required for comprehensive literature reviews, leveraging AI to automate review generation presents a promising opportunity that deserves further exploration. The future development of AI-generated reviews could focus on achieving more precise content-based evaluations of article quality and integrating visual elements to enhance their utility and readability.
Long-term Support Literature Review. The concept of long-term support (LTS) for literature reviews offers a practical solution for maintaining relevance and timeliness in the fast-paced environment of academic research. Unlike traditional reviews, which are generally published once and remain static, an LTS literature review would involve regular updates to include new research findings and advancements in the field. This approach is especially valuable when authors have invested considerable effort in developing structured frameworks or taxonomies that serve as foundational tools in their domain. Such frameworks often require the author’s expertise to adapt effectively to new findings while preserving the original rigor and intent. By maintaining their reviews, authors may ensure that these foundational classifications evolve responsibly, allowing the field to benefit from both continuity and the latest advancements.
@@ 8 Conclusion
In conclusion, this paper conducts a tertiary study on the extensive body of literature reviews in the PAMI field to address three core research questions: (1) What are the structural and statistical features commonly observed in PAMI literature reviews? (2) What strategies can researchers adopt to efficiently navigate the growing number of literature reviews in PAMI? and (3) How can the differences between AI-generated and human-authored reviews be understood to ensure their effective use while mitigating potential risks? To tackle these questions, we perform statistical analysis using the RiPAMI database, and introduce four article-level bibliometric indicators to evaluate review papers from various perspectives.
For the first research question, we conduct an analysis and synthesis of reviews from various disciplines, identifying the dominant structural patterns and shared features specific to PAMI reviews. We develop the RiPAMI database, which contains metadata for over 3,000 reviews. Statistical analysis of this database allows us to pinpoint key statistical features related to word count, reference count, and other metadata elements in PAMI reviews. Furthermore, we utilize the LLM for information extraction, demonstrating that no more than 20% of reviews in the PAMI field comply with PRISMA standards, although this percentage gradually increases each year. We also reveal the positive correlation between the reference quality and academic impact, and examine how the proposed indicators offer insights into the field development trends.
For the second research question, we discuss how the proposed four bibliometric indicators assist researchers in screening review papers. The proposed indicators offer an innovative alternative to traditional bibliometric analysis, enabling a cost-efficient, field-normalized, and real-time assessment of literature reviews. These bibliometric indicators not only furnish clear numerical hints for human researchers but could also offer substantial assistance in the appraisal processes of AI-generated review systems.
For the last research question, this study underscores a significant distinction between human-authored and AI-generated literature reviews: AI-generated reviews typically face challenges in evaluating the potential value of newly published articles. The insights gained from this comparison provide a guide for the effective use of automated review systems and point the way forward for the future development of these systems.
To further support and expand upon this work, we release the code framework that encompasses functionalities including metadata retrieval, indicator calculation, and data analysis, etc. While it is designed for analysis of literature reviews in PAMI, the framework is believed to be universally adaptable, catering to the expansive needs of researchers from disparate fields.
@@
In this section, we provide further insights into the proposed indicators. As presented in Tab. IX, various previous efforts have aimed to tackle the challenge of comparing metrics across fields. For instance, Paper [97] analyzes two well-known field-independent citation metrics at the article level: Field-Weighted Citation Impact (FWCI) and Relative Citation Ratio (RCR). This study suggests that both metrics perform similarly in normalizing citations across research domains. However, these metrics require predefined field categorizations (such as the Scopus All Science Journal Classification) and may not adequately assess papers in emerging subfields. The Field Normalized Citation Success Index (FNCSI), introduced in Refs. [109, 115], is defined as the likelihood that a paper published in Journal A is cited more frequently than a randomly selected paper from Journal B. Although FNCSI offers robustness, it should be noted that it depends on predefined topic keywords and is specifically suited for journal-level assessments.
As the Leiden Manifesto [41] states: metrics should “Account for variation by field in publication and citation practices”. We develop the concept of impact indicators, which is a measure used to gauge the impact of a certain paper in its field. The reason for using the term “indicator” rather than “metric” is we believe that the impact of a certain paper cannot be fully and accurately measured. While metrics like citation counts, h-index, or journal impact factors could indicate a paper’s influence within the academic community, they fail to capture all aspects of its impact. For instance, a research paper might lead to significant advancements in theory, methods, or understanding in its field, none of which would necessarily be reflected in academic metrics. Similarly, a paper might contribute novel concepts or techniques that become influential over time but are not initially evident in citation counts. Hence, the term “indicator” is favored, as it suggests a signal without asserting to encompass the entirety of the academic paper’s contribution.
@@
TABLE IX: Metrics for Evaluating Scholar Impact of Papers: “A” and “J” stand for article-level and journal-level. Field normalized signifies that the metric can be utilized across fields. Value normalized indicates that the range of the value is between 0 and 1.
@@ A.1 TNCSI
$TNCSI$ is a metric used to assess the academic impact of articles across various disciplines. The fundamental distinction of $TNCSI$ from other metrics lies in its independence from the predefined keyword. The requirement for a predefined keyword is a primary reason that most bibliometric indicators face challenges in facilitating cross-disciplinary comparisons. To address such a limitation, we propose adopting ChatGPT [91] (gpt-3.5-turbo-0125) to generate the topic keyword and retrieve related papers according to the keyword. As illustrated in Fig. 9, we ask ChatGPT to identify the most representative topic keyword with the paper’s title and abstract. ChatGPT is one of the most advanced and influential LLMs in the field of natural language processing [149]. Equipped with state-of-the-art language understanding capabilities, ChatGPT has revolutionized the way we interact with AI-powered conversational systems by simply setting “system”, “user”, and “assistant” roles. The “system” role sets the conversation’s behavior and initial context. It provides instructions to guide the assistant’s responses. The “user” role represents the individual interacting with ChatGPT, who inputs messages to the assistant. The “assistant” role is the ChatGPT model itself which would respond based on the provided instructions and user input.
Figure 9: Conceptual illustration of topic keyword generation process. Few-shot prompting may enhance the response quality of the large language model.
@@
where $Citation_{x}$ represents the number of papers with $x$ citations. The $k$ papers related to the topic and their meta-data could be retrieved using a pre-defined topic keyword through online scholar search engines or API, such as Semantic Scholar, CrossRef, or Google Scholar, as mentioned in Sec. 2.3.
@@
In most scenarios, the generated topic word is sufficient to meet expectations, which can be further used as the keyword to retrieve papers from online scholar search engines. Optionally, one can set “System”, “User”, and “Assistant” roles before the final query to improve response quality and create more tailored interactions with the ChatGPT. In other words, a few-shot user-assistant pair prompts the ChatGPT with context on topic granularity. For example, a paper about the classification of irises by an improved CNN may have different perspectives. Some researchers focus more on the algorithm of the improved CNN, while others may be interested in classifying irises. Such ambiguity would likewise make it difficult for ChatGPT to identify the most representative topic keyword as expected. However, this could be addressed by providing the context which consists of (1) the identical prompt template with the replaced title and abstract as user input, and (2) the expected topic keyword as assistant output. By default, we adopt a well-known paper [27] as an example to guide models to generate the topic keyword for all papers.
Similar to other LLMs, ChatGPT adapts its responses based on user-provided natural language prompts. However, natural language prompts can be ambiguous, and different prompts may lead the model to respond with varying quality. Thus, we follow the practices of LLM prompt engineering and carefully design the prompt to optimize the desired output. To determine the optimal prompt, we construct a dataset by manually annotating the topic keywords of 201 papers from various domains published later than the ChatGPT being trained and then compare the performance of multiple prompts on this dataset. The normalized edit distance [141] is adopted to measure the similarity between the GPT-generated keyword and our annotated keyword, where a lower value indicates a higher quality of the prompt. As can be seen from Tab. X, some of the designed prompts achieve decent NED scores for papers in various domains.
@@
3 Identifying the topic of the paper based on the given title and abstract. Avoid using broad or overly general term like ’deep learning’, ’taxonomy’, or ’surveys’. Instead, focus on keyword that is unique and directly pertinent to the paper’s subject. Answer with the word only in the following format: xxx $\times$ 0.36
@@
4 Identifying the topic of the paper based on the given title and abstract. So that I can use it as keyword to search highly related papers from Semantic Scholar. Avoid using broad or overly general term like ’deep learning’, ’taxonomy’, or ’surveys’. Instead, focus on keyword that is unique and directly pertinent to the paper’s subject. Answer with the word only in the following format: xxx $\times$ 0.32
@@
5 Identifying the topic of the paper based on the given title and abstract. I’m going to write a review of the same topic and I will directly use it as keyword to retrieve enough related reference papers in the same topic from scholar search engine. Avoid using broad or overly general term like ’deep learning’, ’taxonomy’, or ’surveys’. Instead, focus on keyword that are unique and directly pertinent to the paper’s subject. Answer with the word only in the following format: xxx $\times$ 0.29
@@
6 Identifying the topic of the paper based on the given title and abstract. I’m going to write a review of the same topic and I will directly use it as keyword to retrieve enough related reference papers in the same topic from scholar search engine. Avoid using broad or overly general term like ’deep learning’, ’taxonomy’, or ’surveys’. Instead, focus on keyword that are unique and directly pertinent to the paper’s subject. Answer with the word only in the following format: xxx $\checkmark$ 0.28
@@
TABLE X: Effectiveness of prompt engineering: The few-shot approach slightly improves the performance of topic keyword extraction.
@@
To provide a clear understanding of the proposed quality indicators, We render the graphical representations of $TNCSI$ in Fig. 10. As can be seen in Fig. 10, $TNCSI$ equals the area under the probability density function curve, which is fitted with the use of maximum likelihood estimation.
Figure 10: Visualization of the proposed $TNCSI$. The $TNCSI$ equals the area under the fitted probability density function curve.
@@ A.2 IEI
To calculate the $IEI$, we first need to obtain the distribution of citation counts over time since publication. Once the citation data is retrieved, we may create a sequence $Seq_{\text{citation}}$ about the number of citations. The $i\in\{0,1,2...,l\}$ item in the sequence $Seq_{\text{citation}}[i]$ represents the number of received citations in the $i_{\text{th}}$ month after the publication, where $l$ represents the number of months allocated for trend observation. Then, a sequence $Seq_{\text{time}}$ of the same length as $Seq_{\text{citation}}$ is generated by enumerating from 0 to $l-1$, Typically, the minimum recommended value for $l$ is 6 or higher. This ensures that the data used for the analysis is adequately representative and the results are reliable. We match the items at the same positions in $Seq_{\text{time}}$ and $Seq_{\text{citation}}$ to determine a set of discrete coordinates $\{(Seq_{\text{time}}[i],Seq_{\text{citation}}[i])\}$, which serve as the control points for shaping the Bézier curve. A Bézier curve is a mathematical representation of smooth curves commonly used in computer graphics, image editing, and design software. The curve starts at the first control point and ends at the last control point, while the intermediate control points influence the curvature and direction of the curve. The number of control points determines the degree $n=l-1$ of the curve.
@@
where $B_{i,n}(t)$ represents the coefficient of the Bézier curve at a given parameter value $t$, which determines the position along the curve ($t=0$ means the start and $t=1$ means the end). $\binom{n}{i}$ is the binomial coefficient, also known as “n choose i”. It represents the number of ways to choose $i$ elements from a set of $n$ elements. $P_{i}$ stands for the $i_{\text{th}}$ control point of the curve.
@@
Finally, the $IEI_{L_{l}}$ can be obtained by averaging the slope of $l=n+1$ distinct points on the curve, as already shown in Eq. (2). Moreover, different months may contribute differently to the $IEI$. For instance, if we desire closer months to have a greater impact, we can achieve this by adjusting the weighting coefficients, $w_{a}$, of the slope at different points to calculate their weighted averages (See in Eq. (12)). In addition, the instantaneous trend could be regarded as the slope of the last month in the sequence. It can be obtained by setting $w_{n}=1$ and the other weighting coefficients to 0. We denote the $IEI$ focused on the last month among the latest $l$ months (excluding the current month) as $IEI_{I_{l}}$, as depicted in Eq. (13).
@@
The illustration of the $IEI$ is shown in Fig. 11. In Fig. 11, the horizontal axis labeled 0 to 5 inversely denotes the months prior to the current month, with 0 representing 6 months ago and 5 denoting the previous month. The corresponding $IEI$ is approximately $-0.76$, which indicates a slightly decreasing citation trend of the investigated review over the last six months.
Figure 11: Visualization of the proposed $IEI$. $IEI$ is the average of the derivatives of each control point on the Bézier curve.

### v6

@@ 6.1 Challenges
Lag and Obsolescence. Scientific knowledge is constantly evolving, yet the process of compiling reviews often struggles to keep up with the latest research advancements—particularly when created by human authors and subject to peer review. Comprehensive reviews are often delayed relative to the emergence of new research areas. For instance, when Mamba [34] started gaining traction, no in-depth reviews were available during its initial months. This delay reflects the time-intensive nature of gathering, evaluating, and synthesizing an ever-expanding body of literature, along with the requirements of the peer-review process.
@@ 6.2 Future
Long-term Support Literature Review. The concept of long-term support (LTS) for literature reviews offers a practical solution for maintaining relevance and timeliness in the fast-paced environment of academic research. Unlike traditional reviews, which are generally published once and remain static, an LTS literature review would involve regular updates to include new research findings and advancements in the field. This approach is especially valuable when structured frameworks or taxonomies have been developed, as they serve as foundational tools in their domain and require continuous refinement to remain useful. While such updates may be curated by the original authors to preserve rigor and intent, automated or AI-assisted mechanisms could also play a critical role in monitoring new publications, suggesting revisions, and integrating fresh evidence. In this way, the field benefits from both the continuity of established classifications and the responsiveness afforded by ongoing, partly automated maintenance.
Intent-Aligned and Customized AI-generated Literature Review. Human-authored reviews are usually designed to be as comprehensive as possible, aiming not only for broad coverage of existing work but also to serve a heterogeneous readership — from newcomers seeking orientation, to experienced researchers looking for synthesis, and practitioners focusing on applications. In contrast, AI-generated reviews are not constrained by this pursuit of exhaustiveness; instead, they can be tailored to align with specific user intentions and deliver the most relevant literature for given purposes. This ability to generate customized reviews highlights AI’s potential to complement traditional human-authored surveys by providing focused, intent-driven perspectives. Future development may further enhance this customization by integrating precise content-based evaluations and visual elements, thereby improving both the relevance and readability of AI-generated reviews.
@@ 7 Conclusion
This study provides the first large-scale tertiary analysis of literature reviews in the PAMI field and yields several key conclusions.
First, our structural investigation of more than 3,000 reviews uncovers highly consistent organizational anchors—such as the placement of introductions and conclusions—yet also reveals systematic variation in and limited adoption of methodological reporting standards. Features like benchmarking, preliminaries, and structured abstracts are gaining momentum, but overall practice remains uneven across subfields and publication types. These findings underscore both the strengths and the blind spots of current review-writing conventions.
Second, our proposed bibliometric indicators ($TNCSI$, RQMRQM, IEIIEI, and $RUI$) demonstrate concrete value in navigating the rapidly growing body of surveys. They help normalize impact across domains, highlight the role of reference quality, track citation momentum, and identify the timeliness of reviews. Together, these measures provide researchers with practical tools for screening overlapping surveys and reducing over-reliance on raw citation counts. They also supply principled signals for future AI-for-Research systems that may incorporate human-authored reviews as seeds for automated synthesis.
Third, our comparative assessment of AI-generated and human-authored reviews points to both progress and limitations. While recent AI systems show improved coherence, organization, and even integration of visual elements, they remain constrained by over-reliance on citation counts, weak personalization, and difficulty in capturing very recent advances. Human experts still excel at nuanced appraisal of new contributions, but AI systems hold promise as intent-driven complements that can deliver customized perspectives for different user needs.
We release our code framework for metadata retrieval, indicator computation, and statistical analysis. Although designed for PAMI, the approach is readily extensible to other domains, and we hope it will serve as a foundation for more reliable, transparent, and purpose-driven review practices across research communities.
Acknowledgements
This research was supported by the Fund of the National Natural Science Foundation of China (Grant No.62576177, 62206134), the Fundamental Research Funds for the Central Universities 070-63233084, and the Tianjin Key Laboratory of Visual Computing and Intelligent Perception (VCIP). Computation is supported by the Supercomputing Center of Nankai University (NKSC). This work was supported by the National Science Fund of China under Grant No. 62361166670. ChatGPT is utilized to refine some of the sentences.
@@ Acknowledgements
In this section, we provide further insights into the proposed indicators. As presented in Tab. 9, various previous efforts have aimed to tackle the challenge of comparing metrics across fields. For instance, Paper [99] analyzes two well-known field-independent citation metrics at the article level: Field-Weighted Citation Impact (FWCI) and Relative Citation Ratio (RCR). This study suggests that both metrics perform similarly in normalizing citations across research domains. However, these metrics require predefined field categorizations (such as the Scopus All Science Journal Classification) and may not adequately assess papers in emerging subfields. The Field Normalized Citation Success Index (FNCSI), introduced in Refs. [110, 117], is defined as the likelihood that a paper published in Journal A is cited more frequently than a randomly selected paper from Journal B. Although FNCSI offers robustness, it should be noted that it depends on predefined topic keywords and is specifically suited for journal-level assessments.
As the Leiden Manifesto [42] states: metrics should ``Account for variation by field in publication and citation practices''. We develop the concept of impact indicators, which is a measure used to gauge the impact of a certain paper in its field. The reason for using the term ``indicator'' rather than ``metric" is we believe that the impact of a certain paper cannot be fully and accurately measured. While metrics like citation counts, h-index, or journal impact factors could indicate a paper's influence within the academic community, they fail to capture all aspects of its impact. For instance, a research paper might lead to significant advancements in theory, methods, or understanding in its field, none of which would necessarily be reflected in academic metrics. Similarly, a paper might contribute novel concepts or techniques that become influential over time but are not initially evident in citation counts. Hence, the term ``indicator'' is favored, as it suggests a signal without asserting to encompass the entirety of the academic paper's contribution.
@@
Table 9: Metrics for Evaluating Scholar Impact of Papers: ``A'' and ``J'' stand for article-level and journal-level. Field normalized signifies that the metric can be utilized across fields. Value normalized indicates that the range of the value is between 0 and 1.
@@ A.1 TNCSI
TNCSITNCSI is a metric used to assess the academic impact of articles across various disciplines. The fundamental distinction of $TNCSI$ from other metrics lies in its independence from the predefined keyword. The requirement for a predefined keyword is a primary reason that most bibliometric indicators face challenges in facilitating cross-disciplinary comparisons. To address such a limitation, we propose adopting ChatGPT [93] (gpt-3.5-turbo-0125) to generate the topic keyword and retrieve related papers according to the keyword. As illustrated in Fig. 12, we ask ChatGPT to identify the most representative topic keyword with the paper's title and abstract. ChatGPT is one of the most advanced and influential LLMs in the field of natural language processing [157]. Equipped with state-of-the-art language understanding capabilities, ChatGPT has revolutionized the way we interact with AI-powered conversational systems by simply setting ``system'', ``user'', and ``assistant'' roles. The ``system'' role sets the conversation's behavior and initial context. It provides instructions to guide the assistant's responses. The ``user'' role represents the individual interacting with ChatGPT, who inputs messages to the assistant. The ``assistant'' role is the ChatGPT model itself which would respond based on the provided instructions and user input.
Figure 12: Conceptual Illustration of Topic Keyword Generation Process: Few-shot prompting may enhance the response quality of the large language model.
@@
where $Citation_{x}$ represents the number of papers with $x$ citations. The $k$ papers related to the topic and their meta-data could be retrieved using a pre-defined topic keyword through online scholar search engines or API, such as Semantic Scholar, CrossRef, or Google Scholar.
@@
In most scenarios, the generated topic word is sufficient to meet expectations, which can be further used as the keyword to retrieve papers from online scholarly search engines. Optionally, one can set ``System'', ``User'', and ``Assistant'' roles before the final query to improve response quality and create more tailored interactions with the ChatGPT. In other words, a few-shot user-assistant pair prompts the ChatGPT with context on topic granularity. For example, a paper about the classification of irises by an improved CNN may have different perspectives. Some researchers focus more on the algorithm of the improved CNN, while others may be interested in classifying irises. Such ambiguity would likewise make it difficult for ChatGPT to identify the most representative topic keyword as expected. However, this could be addressed by providing the context which consists of (1) the identical prompt template with the replaced title and abstract as user input, and (2) the expected topic keyword as assistant output. By default, we adopt a well-known paper [28] as an example to guide models to generate the topic keyword for all papers.
Similar to other LLMs, ChatGPT adapts its responses based on user-provided natural language prompts. However, natural language prompts can be ambiguous, and different prompts may lead the model to respond with varying quality. Thus, we follow the practices of LLM prompt engineering and carefully design the prompt to optimize the desired output. To determine the optimal prompt, we construct a dataset by manually annotating the topic keywords of 201 papers from various domains published later than the ChatGPT being trained and then compare the performance of multiple prompts on this dataset. The normalized edit distance [148] is adopted to measure the similarity between the GPT-generated keyword and our annotated keyword, where a lower value indicates a higher quality of the prompt. As can be seen from Tab. 10, some of the designed prompts achieve decent NED scores for papers in various domains.
@@
3 Identifying the topic of the paper based on the given title and abstract. Avoid using broad or overly general term like 'deep learning', 'taxonomy', or 'surveys'. Instead, focus on keyword that is unique and directly pertinent to the paper's subject. Answer with the word only in the following format: xxx $\times$ 0.36
@@
4 Identifying the topic of the paper based on the given title and abstract. So that I can use it as keyword to search highly related papers from Semantic Scholar. Avoid using broad or overly general term like 'deep learning', 'taxonomy', or 'surveys'. Instead, focus on keyword that is unique and directly pertinent to the paper's subject. Answer with the word only in the following format: xxx $\times$ 0.32
@@
5 Identifying the topic of the paper based on the given title and abstract. I'm going to write a review of the same topic and I will directly use it as keyword to retrieve enough related reference papers in the same topic from scholar search engine. Avoid using broad or overly general term like 'deep learning', 'taxonomy', or 'surveys'. Instead, focus on keyword that are unique and directly pertinent to the paper's subject. Answer with the word only in the following format: xxx $\times$ 0.29
@@
6 Identifying the topic of the paper based on the given title and abstract. I'm going to write a review of the same topic and I will directly use it as keyword to retrieve enough related reference papers in the same topic from scholar search engine. Avoid using broad or overly general term like 'deep learning', 'taxonomy', or 'surveys'. Instead, focus on keyword that are unique and directly pertinent to the paper's subject. Answer with the word only in the following format: xxx $\checkmark$ 0.28
@@
Table 10: Effectiveness of Prompt Engineering: The few-shot approach slightly improves the performance of topic keyword extraction.
It is noteworthy that the proposed indicator exhibits a certain degree of robustness with respect to the choice of keywords. This robustness stems from the underlying mechanisms of modern search engines, which rely primarily on semantic similarity rather than strict keyword matching. As a result, even when the keywords differ, the retrieved literature remains largely consistent. In Table 11, we report the KL divergences of citation distributions obtained from searches with different but semantically related keywords. The results show that the average KL divergence remains below 0.1, indicating that the metric is highly stable under variations in keyword phrasing.
Anchor Term Comparison Terms Avg. KL
Object Detection Target Detection, Object Localization 0.080
Face Detection Facial Landmark Detection, Face Recognition 0.052
Image Classification Object Categorization, Visual Classification 0.086
Pose Estimation Human Pose Detection, Body Pose Estimation 0.194
Semantic Segmentation Scene Segmentation, Class-Level Segmentation 0.049
Generative Adversarial Networks GAN, Adversarial Networks 0.014
Object Tracking Visual Tracking, Moving Object Detection 0.075
Image Super-Resolution Image Enhancement, Super-Resolution Imaging 0.014
Action Recognition Activity Recognition, Gesture Recognition 0.090
3D Reconstruction Three-Dimensional Reconstruction, Scene Reconstruction 0.010
Pedestrian Detection Person Localization, Human Detection 0.042
Vehicle Detection Car Detection, Automobile Detection 0.113
Text-to-Image Generation Visual Synthesis, Image Generation from Text 0.020
Neural Style Transfer Style Transfer, Image Transformation 0.147
Speech Recognition Voice Recognition, Speech-to-Text 0.044
Object Recognition Object Identification, Visual Recognition 0.076
Medical Imaging Radiology Imaging, Clinical Imaging 0.011
Robotic Vision Machine Vision, Automated Vision 0.022
Data Augmentation Data Synthesis, Synthetic Data Generation 0.179
Human-Computer Interaction User Interface Interaction, HCI 0.005
Scene Graph Generation Object-Relationship Graph, Contextual Graphs 0.161
Overall 0.071
Table 11: Average KL Divergence of Synonym Groups: The metric is computed in a single direction with the first term of each group as the anchor, and the ``Comparison Terms'' column lists the alternative expressions. These values are generally low, indicating robustness under keyword variations.
@@
To provide a clear understanding of the proposed quality indicators, we render the graphical representations of $TNCSI$ in Fig. 13 left panel. As can be seen in Fig. 13, $TNCSI$ equals the area under the probability density function curve, which is fitted with the use of maximum likelihood estimation. In the right panel, $TNCSI$ is contrasted with the percentile-based citation indicator, which corresponds to the empirical cumulative distribution function (ECDF). Unlike the smooth curve of $TNCSI$, the ECDF is discrete, meaning that different citation counts may share the same percentile value. This discreteness can introduce unavoidable errors in certain tasks such as network training [156].
Figure 13: Illustration of the Proposed Quality Indicators: The $TNCSI$ is defined as the area under the fitted probability density function curve estimated via maximum likelihood, resulting in a smooth cumulative representation. In contrast, the percentile-based citation indicator corresponds to the empirical cumulative distribution function (ECDF), which is discrete and may assign the same percentile value to different citation counts.
@@ A.2 IEI
To calculate the $IEI$, we first need to obtain the distribution of citation counts over time since publication. Once the citation data is retrieved, we may create a sequence $Seq_{\text{citation}}$ about the number of citations. The $i\in\{0,1,2...,l\}$ item in the sequence $Seq_{\text{citation}}[i]$ represents the number of received citations in the $i_{\text{th}}$ month after the publication, where $l$ represents the number of months allocated for trend observation. Then, a sequence $Seq_{\text{time}}$ of the same length as $Seq_{\text{citation}}$ is generated by enumerating from 0 to l1l-1, Typically, the minimum recommended value for $l$ is 6 or higher. This ensures that the data used for the analysis is adequately representative and the results are reliable. We match the items at the same positions in $Seq_{\text{time}}$ and $Seq_{\text{citation}}$ to determine a set of discrete coordinates $\{(Seq_{\text{time}}[i],Seq_{\text{citation}}[i])\}$, which serve as the control points for shaping the Bézier curve. A Bézier curve is a mathematical representation of smooth curves commonly used in computer graphics, image editing, and design software. The curve starts at the first control point and ends at the last control point, while the intermediate control points influence the curvature and direction of the curve. The number of control points determines the degree n=l1n=l-1 of the curve.
@@
where $B_{i,n}(t)$ represents the coefficient of the Bézier curve at a given parameter value $t$, which determines the position along the curve ($t=0$ means the start and $t=1$ means the end). $\binom{n}{i}$ is the binomial coefficient, also known as ``n choose i". It represents the number of ways to choose $i$ elements from a set of $n$ elements. $P_{i}$ stands for the $i_{\text{th}}$ control point of the curve.
@@
Finally, the $IEI_{L_{l}}$ can be obtained by averaging the slope of $l=n+1$ distinct points on the curve. Moreover, different months may contribute differently to the $IEI$. For instance, if we desire closer months to have a greater impact, we can achieve this by adjusting the weighting coefficients, $w_{a}$, of the slope at different points to calculate their weighted averages (See in Eq. (12)). In addition, the instantaneous trend could be regarded as the slope of the last month in the sequence. It can be obtained by setting $w_{n}=1$ and the other weighting coefficients to 0. We denote the $IEI$ focused on the last month among the latest $l$ months (excluding the current month) as $IEI_{I_{l}}$, as depicted in Eq. (13).
@@
The illustration of the $IEI$ is shown in Fig. 14. In Fig. 14, the horizontal axis labeled 0 to 5 inversely denotes the months prior to the current month, with 0 representing 6 months ago and 5 denoting the previous month. From the figure it can be observed that, compared with the fifth-degree polynomial fitting, the $IEI$ avoids overfitting and produces a smoother and more stable trend. In contrast to the first- or third-degree polynomial fittings, which either oversimplify the dynamics or introduce spurious inflection points, the $IEI$ achieves a balanced representation that captures the overall trajectory while maintaining robustness against local noise.
Figure 14: Visualization of the Proposed IEIIEI: IEIIEI is the average of the derivatives of each control point on the Bézier curve.